# Exploring the Latent Capacity of LLMs for One-Step Text Generation thanks: Under review

### v1

@@ Abstract
autoregressive generation from just one specially trained input embedding. In this work, we explore whether such reconstruction is possible without autoregression.We show that frozen LLMs can generate hundreds of accurate tokens in just one forward pass, when provided with only two learned embeddings. This reveals a surprising and underexplored capability of LLMs – multi-token generation without iterative decoding. We investigate the behaviour of these embeddings and provide insight into the type of information they encode. We also empirically show that although these representations are not unique for a given text, they form connected and local regions in embedding space – a property that suggests the potential of learning a dedicated encoder into that space.
Exploring the Latent Capacity of LLMs for One-Step Text Generation thanks: Under review
@@ 1 Introduction
Large language models (LLMs) are typically trained to generate text in an autoregressive manner – they predict one token at a time based on the previously generated context.
Recent work by Kuratov et al. (2025) demonstrated that LLMs can autoregressively generate an arbitrary text starting from a single, specially trained input embedding corresponding to that text. This raises an intriguing question: is autoregressive generation an essential part of such reconstruction? Or, in other words, can LLMs reconstruct accurate multi-token sequences from some compressed representation in a single forward pass, without any iterative generation, and if so, how?
In this work, we show that this is possible, investigate what those compressed representations encode and whether this finding reveals anything about LLMs’ parallel generation capabilities.
@@
1. We show that LLMs can reconstruct arbitrary sequences from as few as two learned input embeddings, achieving perfect reconstruction of sequences of up to several hundred tokens.
@@
Figure 1: Two "proto-tokens" (trainable embeddings) are fed into frozen, pre-trained LLM and optimized in such a way, that the LLM predicts an arbitrary target token-sequence in a single forward pass. One of the "proto-tokens" ($e_{t}$) is trained for each text separately, while the other ($m$) could be reused.
@@ 2 Related Work
The most direct influence for our work is a paper by Kuratov et al. (2025), which showed that frozen LLMs can reconstruct an arbitrary sequence of tokens $T=[t_{1},\dots,t_{N}]$ if given a set of special, so-called memory tokens $[mem_{1},\dots,mem_{K}]$. The embeddings for these tokens are trained by optimizing a causal language modeling objective (next-token prediction cross-entropy loss) over a concatenated input sequence $Z=[mem_{1},\dots,mem_{K},t_{1},\dots,t_{N}]$ passed through a frozen LLM. In the case of perfect next-token prediction accuracy (which could be achieved for reasonable text length), this allows the model to autoregressively predict the whole text starting from the memory tokens. The number of memory tokens controls the maximum text length and can be as low as one.
Although surprisingly long (up to 1568 tokens) texts could be compressed even into a single memory token, the authors note that the embeddings trained from different random initializations for the same text often end up far apart. Moreover, linear interpolations between those embeddings produce very poor reconstruction accuracy, suggesting that the solution space lacks desirable smoothness and locality qualities, which are important for learning a practical encoder that could replace the direct optimization.
Our work also relates to efforts in prompt-tuning and its variants (Lester et al., 2021; Liu et al., 2024; Li and Liang, 2021). Most similarly, Lester et al. (2021) train task-specific soft tokens to condition the frozen LLMs to improve their performance in new tasks. Finally, several speculative (Xia et al., 2023) and parallel (Santilli et al., 2023) decoding approaches utilize a similar mechanism for multiple token prediction using decoder architectures. More specifically, they add special [PAD] or [MASK] tokens at the end of the current context in order to make a prediction for several tokens into the future at once. Critically, in these works either special training or multiple generative iterations are required.
@@ 3 Method
To adopt the approach from Kuratov et al. (2025) to a non-autoregressive case, we replace all input tokens of the LLM with specially trained "proto-tokens" and predict the target token sequence in one forward pass. In practice, "proto-tokens" are just trainable vectors that are not tied to any real items in the vocabulary. The main difference between regular tokens and these "proto-tokens" is that "proto-tokens" encode multiple tokens at once and only produce human-readable text after passing through the LLM. Our goal is to identify the smallest possible number of such "proto-tokens" needed for accurate reconstruction. Interestingly, we find that it is essential to have at least two – the performance drops dramatically when using only one (see Section 4).
There are many ways to arrange two vectors as an input sequence of arbitrary length. We report results for different variants later in the paper, but here we describe the arrangement that is used in the majority of the experiments.
@@
$Z=[e,m,m,\dots,m]$ – one copy of token $e$ is followed by $N-1$ copies of token $m$, where $N$ is the target text length. We then train the vectors by optimizing cross-entropy loss between the target sequence $T=[t_{1},t_{2},\dots,t_{N}]$ and the frozen LLM’s output for the input sequence. The prediction is obtained using standard causal attention masking, so that the predicted probabilities for the token $t_{i}$ depend on the first $i$ input "proto-tokens" (see Figure 1).
@@
To gain insights into the structure of the solution space of our problem, we analyze whether different proto-token embeddings obtained for the same text but from different random initializations are connected. We adopt a technique from (Garipov et al., 2018) which is used to find paths connecting different minima of the loss function in computer vision tasks. We optimize the parameters of a degree-one Bezier curve, connecting two solutions, to maximize reconstruction accuracy along the curve. The curve is parameterized by a control point $\pi$ in the following way:
$\phi_{\pi}(t)=(1-t)^{2}p_{1}+2t(1-t)\pi+t^{2}p_{2}$ (3)
@@
The expectation of the cross-entropy loss function under the uniform distribution over $t\in[0,1]$ (4) is minimized by iteratively sampling $\tilde{t}\in[0,1]$ and making a gradient step, effectively obtaining unbiased estimate of the gradient of $l_{\pi}$:
$l_{\pi}=\int_{0}^{1}{\sum_{i=1}^{N}-log\mathbb{P}_{LM}(t_{i}|\phi_{\pi}(t))dt}$ (4)
This acts as a more tractable alternative to direct optimization under the uniform distribution along the curve itself.
@@ 4 Experiments and results
The embeddings of the proto-tokens are initialized randomly from a standard normal distribution and optimized using AdamW optimizer (Loshchilov and Hutter, ) with $0.01$ learning rate, $\beta_{1}$, $\beta_{2}$ set to $0.9$ and a weight decay of $0.01$. The embeddings are trained for $5000$ iterations with early stopping if perfect reconstruction accuracy is achieved. This number of iterations is often insufficient for convergence, but due to limited computational resources, we are unable to increase it. Instead, we aggregate results across multiple sequences. All models are loaded and trained using PyTorch framework and the Hugging Face Transformers library. Each experimental run is done on a single A100 or H100 80GB GPU with gradient accumulation enabled where necessary.
The code is available at this page [^3] https://github.com/Glebzok/OneStepLLMGeneration.
@@
To select the best way to arrange two proto-tokens as input to an LLM for the main experiments, we conduct test runs on a single dataset-model pair for the variety of arrangements. For each arrangement, the same 50 texts from the PG-19 are selected, and the Llama-3.2-1B model is trained on prefixes of these texts at lengths [1, 2, 4, 8, 16, 32, 64, 128, 256, 512, 1024] to assess how token-level reconstruction accuracy changes with respect to sequence length $N$. A representative selection of results is presented in Table 1.
@@
[…]
Table 1: Reconstruction accuracies for different input token arrangements across varying sequence lengths. Subscripts indicate the number of copies for each proto-token. The last two schemes differ as follows: with the first one, the LLM is trained to predict the first text token $t_{1}$ for the proto-token $e$, while with the second one, the prediction for proto-token $e$ is not guided and $t_{1}$ is a target prediction for the first copy of $m$ instead.
Interestingly, having two trainable tokens is essential for the performance – the scheme with one trainable token fails to reconstruct even 2-token text, while best two-token schemes can reconstruct $256$-token texts almost perfectly.
Moreover, the way these two tokens are arranged is also important, with the best results obtained when the first token $e$ is followed by $N-1$ copies of the second token $m$. This asymmetrical arrangement and critical necessity for two tokens suggest possible variation in functions of $e$ and $m$. It is possible, that while one of them mostly incorporates language information, the role of the other one is mainly structural or mechanistic. This could be related to the phenomenon of "attention sinks" – Xiao et al. (2023) showed that LLMs strongly attend to the initial tokens in the sequence even when they are not relevant. Moreover, adding a placeholder token as an attention sink could largely improve the performance of window-attention based models, which do not see the initial tokens by design. So, it is possible, that in order to successfully decode "information" proto-token, LLM needs a distinguishable "sink" proto-token, which can be used as attention sink.
@@
In the previous section, we showed that the quality of reconstruction is very dependent on having two separate proto-tokens as an input. This observation, led us to hypothesize that, a second token plays some structural or mechanistic purposes and does not contain information about the sequence itself. In that case, the second token could be shared between texts, reducing the number of optimized parameters, and simplifying the training process of the potential encoder.
To test this hypothesis, we run the same optimization process, but splitting $256$ texts from the PG-19 dataset into groups of different sizes $S_{g}\in[1,4,16,64,256]$ and sharing either $e$ or $m$ within each group. We selected the maximum length of the text that can be losslessly compressed in a non-sharing mode - $256$. The results are averaged over $10$ random seeds. The selection of the results is presented in Table 2.
@@
Table 2: Reconstruction accuracy for schemes where one of the trainable tokens is shared within a group across different group sizes. "max" aggregation indicates that for every text, maximum accuracy across ten random seeds is selected and then averaged across texts, while "avg" denotes averaging across both seeds and texts.
Sharing either token yields comparable performance if provided with sufficiently large number of restarts (random seeds), but the required number of restarts increases significantly with group size.
Depending on the proto-token being shared, we can build different intuitions behind the function of the shared tokens and the method itself. If $e$-token is shared, which is located in the very beginning of the input sequence, the analogy that comes to mind is prompt-tuning Lester et al. (2021), where a set of prompt embeddings is trained in order to improve performance in some specific task. In our case, a shared token $e$ could be viewed as an "instruction" saying what an LLM should do with the upcoming embeddings ($m$-tokens) – decode different pieces of information for different positions. If $m$ is shared, then training and prediction scheme resembles some of the speculative decoding approaches Xia et al. (2023), where a number of special "[mask]" tokens are appended at the end of the sequence and the prediction for all them is then done in parallel. For all other experiments, unless stated otherwise, we use scheme with sharing $m$ token between texts and random seeds and $e$ token being unique for each text/seed pair.
Share $p$ Pythia Llama
@@
We already see, that similar to autoregressive mode (Kuratov et al., 2025), LLMs can generate fairly long sequences in just one forward pass. To characterize this capability, and understand how it scales with model size, we run the optimization process for text prefixes of the predefined lengths [4, 5, 8, 11, 16, 22, 32, 45, 64, 90, 128, 181, 256, 362, 512, 724, 1024, 1448]. We report the maximum values of $C_{tokens}$, and $H_{max}$ which correspond to the longest prefix for which at least $0.99$ token-level accuracy is achieved – we treat such sequences as successfully predicted. In addition to a scheme with a shared $p$ token, we also run a scheme with $p$ not shared, to eliminate the effect of the insufficient number of random initializations. While our results in Section 4, suggest that $p$, can in principal, be shared without any quality drop, we also note that the optimization process is highly sensitive to initialization, especially when the proto-tokens are shared. The results are presented in Table 3.
@@
Larger models in Llama family show greater reconstruction capabilities than the smaller ones of their family, while the situation with Pythia model-family is less obvious, with all the models showing approximately the same performance. Llama 1B model is also able to reconstruct almost three times larger sequence compared to Pythia model of the same size.
The source of the natural language (unseen / seen / generated) doesn’t seem to have any systematic influence on the quality of reconstruction in terms of the number of tokens, while for unnatural random texts the generation capacity is significantly worse. This suggests that our "proto-tokens" do not "store" text tokens directly, but encode some more high-level representations, using language modeling capabilities of LLM. However, we also can’t say that the compressibility of the text is determined by its likelihood under the sequential language model. In fact, we observe the opposite trend: lower total information content $H_{LM}$ is compressed for less-information dense texts, such as generated by the LLM itself.
@@
Figure 2: Maximum language information ($H_{LM}$ for a maximum text prefix that is accurately reconstructed) compressed for different models and datasets. In the left plot, a single [mem] token is used in the autoregressive setting, and in the non-autoregressive one, $p$ proto-token is shared between all texts within each model. In the right plot, two [mem] tokens are used and $p$ proto-tokens are not shared. Each small point on the plots represents a single text, larger points indicate the average within each (model, dataset) pair.
@@
This difference is highlighted in Figure 2, where the amount of the language-information contained in trainable tokens is compared to autoregressive setup. The performance for unnatural texts is very similar and sometimes even identical, while for natural texts, the difference in capacity can be up to five times lower. However, more often the performance is just two times lower in non-autoregressive case, suggesting that autoregressive decoding approximately doubles the "effective" information density for natural texts – the density of the information that could be effectively decoded.
@@
Figure 3: Reconstruction throughput comparison between autoregressive and non-autoregressive setups. For each (model, dataset) pair, the throughput is calculated as a maximum losslessly compressible length divided by the reconstruction time. To measure reconstruction time, we use PyTorch profiling tools.
@@
Although less information-dense, our one-forward method achieves significantly higher decoding throughput in the context of text reconstruction – outperforming its autoregressive counterpart by a factor of 279 on average (Figure 3). This dramatic difference is primary due to the number of forward passes. While an obvious downstream task is still to be found, such speed could matter for fast context-compression and decompression, on-device inference, or a setting where decoding speed is particularly important.
@@
We start by measuring the distances between three types proto-token embedding pairs: 1) corresponding to the same generated sequence, but different random seeds, 2) corresponding to the different texts but generated from the same context, 3) corresponding to the different texts generated from different contexts. As shown in Figure 4, the same-text solutions are almost always located closer to each other than different-texts solutions, which suggests locality in the learned representations. At the same time, same-context solutions are noticeably closer to each other than different-context ones. This may indicate that the encoded information at least partially reflects the potential context of the text. However, we should be careful to account for the texts generated from the same context being more similar in general.
To do that, we measure pairwise distances between generated texts, and examine whether the distance between learned proto-token embeddings differ for a fixed distance between the texts. We use token-level measure of text similarity and semantic-level measure (see Section 3). For both measures, (Figure 5) we observer that, given similar distances between texts, the proto-token embeddings are consistently closer when the texts originate from the same context. We conclude that learned proto-tokens contain information beyond the information about the target sequence itself – it somehow describes the potential context of the sequence.
Kuratov et al. (2025) raised the following concern about the structure of the solution space in autoregressive setup. Even though the same-text token embeddings are on average closer to each other than different-text token embeddings, they seem to be disconnected – a linear interpolation between two solutions does not yield a valid reconstruction. This could mean that the potential encoding to this space could be problematic as the same object could be mapped to disconnected regions. We find that in our non-autoregressive case, the linear interpolation between same-text solutions also does not produce a solution (Figure 6).
Figure 6: Pairwise interpolation accuracies between $10$ solutions for $5$ texts ($5\times 10\times 9/2$ pairs in total).

### v2

# Exploring the Hidden Capacity of LLMs for One-Step Text Generation

@@ Abstract
autoregressive generation from just one trained input embedding. In this work, we explore whether autoregressive decoding is essential for such reconstruction. We show that frozen LLMs can generate hundreds of accurate tokens in just one token-parallel forward pass, when provided with only two learned embeddings. This reveals a surprising and underexplored multi-token generation capability of autoregressive LLMs. We examine these embeddings and characterize the information they encode. We also empirically show that, although these representations are not unique for a given text, they form connected and local regions in embedding space – suggesting the potential to train a practical encoder. The existence of such representations hints that multi-token generation may be natively accessible in off-the-shelf LLMs via a learned input encoder, eliminating heavy retraining and helping to overcome the fundamental bottleneck of autoregressive decoding while reusing already-trained models.
Exploring the Hidden Capacity of LLMs for One-Step Text Generation
@@ 1 Introduction
Large language models are typically trained to generate text in an autoregressive manner – they predict one token at a time based on the previously generated context. Several attempts aim to change this. However, they either require an additional model for candidate generation (Leviathan et al., 2023), substantial additional fine-tuning of autoregressive LLM (Cai et al., 2024; Stern et al., 2018; Gloeckle et al., 2024) or full model retraining (Ghazvininejad et al., 2019; Austin et al., 2021; Li et al., 2022). This leaves an open question – is it possible to reuse autoregressively pretrained LLM for multi-token generation with minimal to no additional training. We discover a previously undocumented phenomenon, that can help us to achieve this goal.
We found that for any given text of reasonable length, there exists a latent one-vector representation of this text, such that, if a frozen pretrained LLM is conditioned on this representation, it accurately generates the whole text in a single forward pass, without any iterative decoding. In this work, we demonstrate this phenomenon, investigate what those compressed representations encode and whether this finding reveals anything about LLMs’ parallel generation capabilities.
Figure 1: One pass, many tokens. Each dot shows the maximum exact reconstruction length in a single non-autoregressive forward pass with frozen weights, conditioned only on two learned embeddings – evidence of hidden multi-token capabilities.
@@
1. We show that LLMs can reconstruct arbitrary sequences of hundreds of tokens from as few as two learned input embeddings, with one of them being universal for all texts.
@@
The code is available at this GitHub page.
@@ 2 Related Work
The most direct influence for our work is a paper by Kuratov et al. (2025), which showed that frozen LLM can reconstruct an arbitrary sequence of tokens $T=[t_{1},\dots,t_{N}]$ if given a sequence of special, so-called memory tokens $mem_{1},\dots,mem_{K}]$. The embeddings for these tokens are trained by optimizing a causal language modeling objective over a concatenated input sequence $Z=[mem_{1},\dots,mem_{K},t_{1},\dots,t_{N}]$} passed through a frozen LLM. In the case of perfect next-token prediction accuracy (which could be achieved for reasonable text length), this allows the model to autoregressively predict the whole text starting from the memory tokens. The number of memory tokens controls the maximum text length and can be as low as one.
Although surprisingly long (up to 1568 tokens) texts could be compressed even into a single memory token, the authors note that the embeddings trained from different random initializations for the same text end up far apart. Moreover, linear interpolations between those embeddings produce poor reconstruction accuracy, suggesting that the solution space lacks desirable smoothness and locality qualities, which are important for learning a practical encoder that could replace direct optimization.
Our work also relates to efforts in prompt-tuning and its variants (Lester et al., 2021; Liu et al., 2024; Li and Liang, 2021). Most similarly, Lester et al. (2021) train task-specific soft tokens to condition frozen LLMs to improve their performance on new tasks. Several speculative (Xia et al., 2023) and parallel (Santilli et al., 2023) decoding approaches utilize a similar mechanism for multiple token prediction using decoder architectures. More specifically, they add special [PAD] or [MASK] tokens at the end of the current context in order to make a prediction for several tokens into the future at once. Critically, in these works either special training or multiple generative iterations are required.
@@ 3 Method
To adopt the approach from Kuratov et al. (2025) to the non-autoregressive case, we replace all input tokens of the LLM with specially trained "proto-tokens" and predict the target token sequence in one forward pass. In practice, "proto-tokens" are just trainable vectors that are not tied to any real items in the vocabulary. The main difference between regular tokens and these "proto-tokens" is that "proto-tokens" encode multiple tokens at once and only produce human-readable text after passing through the LLM. Our goal is to identify the smallest possible number of such "proto-tokens" needed for accurate reconstruction. Interestingly, we find that it is essential to have at least two – the performance drops dramatically when using only one (see Section 4).
There are many ways to arrange two vectors into an input sequence of arbitrary length. We report results for different variants later in the paper, but here we describe the arrangement that is used in the majority of the experiments.
@@
Figure 2: Two "proto-tokens" (trainable embeddings) are fed into frozen, pretrained LLM and optimized in such a way that LLM predicts an arbitrary token-sequence in a single forward pass. $e_{t}$ is trained for each text separately, while $m$ could be shared across texts.
@@
$Z=[e,m,m,\dots,m]$ – one copy of token $e$ is followed by N1N-1 copies of token $m$, where $N$ is the target text length. We then train the vectors by optimizing cross-entropy loss between the target sequence $T=[t_{1},t_{2},\dots,t_{N}]$ and the frozen LLM’s output for the input sequence:
$L_{CE}=-\sum_{i=1}^{N}log\mathbb{P}_{LM}(t_{i}\mid e,\underbracket{m,\dots,m}_{i-1})$ (1)
The prediction is obtained using standard causal attention masking, so that the predicted probabilities for the token $t_{i}$ depend on the first $i$ input "proto-tokens" (see Figure 2).
@@
To gain insights into the structure of the solution space of our problem, we analyze whether different proto-token embeddings obtained for the same text but from different random initializations are connected. We adopt a technique from (Garipov et al., 2018) which is used to find paths connecting different minima of the loss function in computer vision tasks. We optimize the parameters of a degree-two Bezier curve, connecting two solutions, to maximize reconstruction accuracy along the curve. The curve is parameterized by a control point $\pi$ in the following way:
$\phi_{\pi}(\tau)=(1-\tau)^{2}p_{1}+2\tau(1-\tau)\pi+\tau^{2}p_{2}$ (4)
@@
We want to find the value of $\pi$ that minimizes the cross-entropy loss along the curve. To do that, we minimize the expectation of the cross-entropy loss with respect to a uniform distribution of $\tau$:
$l_{\pi}=\mathop{\mathbb{E}}_{\tau\sim\mathcal{U}[0,1]}{\sum_{i=1}^{N}-log\mathbb{P}_{LM}(t_{i}\mid\phi_{\pi}(\tau))}$ (5)
To do that, we iteratively sample $\tau\sim\mathcal{U}[0,1]$ and optimize $l_{\pi}$ with respect to $\pi$ using Adam optimizer. This optimization under the uniform distribution over $\tau$ acts as a more tractable alternative to direct optimization under the uniform distribution along the curve itself.
@@ 4 Experiments and results
The embeddings of the proto-tokens are initialized from standard normal distribution and optimized via AdamW optimizer (Loshchilov and Hutter, ) with $0.01$ learning rate, $\beta_{1}$, $\beta_{2}$ set to $0.9$ and a weight decay of $0.01$. The embeddings are trained for $5000$ iterations with an early stopping at a perfect reconstruction accuracy. This number of iterations is often insufficient for convergence, but due to limited computational resources, we are unable to increase it. Instead, we aggregate results across multiple sequences and report the best results. Although, the exact reconstruction capacity could be under-estimated, we believe that, given he exploratory nature of this work, it is more important to demonstrate and characterize the phenomenon itself, rather than to achieve the precise upper bound on reconstruction capacity. All models are trained using PyTorch framework and Transformers library. Each experimental run is done on a single A100 or H100 80GB GPU with gradient accumulation enabled where necessary.
@@
To select the best way to arrange two proto-tokens as input for LLM for the main experiments, we conduct test runs on a single dataset-model pair for the variety of arrangements. For each arrangement, the same 50 texts from the PG-19 are selected, and the Llama-3.2-1B model is trained on prefixes of these texts at lengths [1, 2, 4, 8, 16, 32, 64, 128, 256, 512, 1024] to assess token-level reconstruction accuracy change with respect to sequence length $N$. A representative selection of results is presented in Table 1.
@@
$\begin{aligned} {}_{{}_{\times N}}\end{aligned}$ $1.00_{{}_{\pm 0.00}}$ $0.45_{{}_{\pm 0.31}}$ $17_{{}_{\pm 0.18}} 0.0$$01_{{}_{\pm 0.01}}$
$\begin{aligned} {}_{{}_{\times(N/2)}}[m]_{{}_{\times(N/2)}}\end{aligned}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $12_{{}_{\pm 0.13}} 0.0$$01_{{}_{\pm 0.01}}$
$\begin{aligned} {}_{{}_{\times(N/2)}}\end{aligned}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $0.17_{{}_{\pm 0.34}}$
$\begin{aligned} [m]_{{}_{\times N}}\end{aligned}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $0.97_{{}_{\pm 0.15}}$
$\begin{aligned} [m]_{{}_{\times(N-1)}}\end{aligned}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $1.00_{{}_{\pm 0.00}}$ $0.99_{{}_{\pm 0.10}}$
Table 1: Reconstruction accuracies for different input token arrangements across varying sequence lengths. Subscripts indicate the number of copies for each proto-token. In the second-to-last scheme the LLM is trained to predict the first text token $t_{1}$ for the proto-token $e$, while with the last one, the prediction for proto-token $e$ is not guided and $t_{1}$ is a target for the first copy of $m$.
Interestingly, having two proto-tokens is essential. The one-token scheme fails to reconstruct even 2-token text, while best two-token schemes reconstruct $256$-token texts almost perfectly.
Moreover, the way these two tokens are arranged is also important, with the best results obtained when the first token $e$ is followed by N1N-1 copies of the second token $m$. This asymmetrical arrangement and critical necessity for two tokens suggest possible variation in functions of $e$ and $m$. It is possible, that while one of them mostly incorporates language information, the role of the other one is mainly structural or mechanistic. This could be related to the phenomenon of "attention sinks" – Xiao et al. (2023) showed that LLMs strongly attend to the initial tokens in the sequence even when they are not relevant. So, it is possible, that in order to successfully decode "information" proto-token, LLM needs a distinguishable "sink" proto-token, which can be used as attention sink.
@@
In the previous section, we showed that the quality of reconstruction is very dependent on having two separate proto-tokens as an input. This observation led us to hypothesize that, a second token serves a structural or mechanistic purpose and does not contain information about the sequence itself. In that case, the second token could be shared between texts, reducing the number of optimized parameters, and simplifying the training process of the potential encoder.
To test this hypothesis, we run the same optimization process, splitting $256$ texts from the PG-19 into groups of varying sizes $S_{g}\in[1,4,16,64,256]$ and sharing either $e$ or $m$ within each group. We selected the maximum length of the text that can be losslessly compressed in a non-shared mode - $256$. The results are averaged over $10$ random seeds. The selection of results is presented in Table 2.
@@
Table 2: Reconstruction accuracy with one of proto-tokens shared within groups for different group sizes. "max" indicates that for every text, maximum accuracy across ten random seeds is averaged across texts, while "avg" denotes averaging across both seeds and texts.
Sharing either token yields comparable performance if provided with a sufficiently large number of restarts (random seeds), but the required number of restarts increases significantly with group size.
Share $m$ Pythia Llama
@@
Depending on the proto-token being shared, we can build different intuitions behind the function of the shared tokens and the method itself.
If the $e$-token is shared, which is located in the very beginning of the input sequence, the analogy that comes to mind is prompt-tuning Lester et al. (2021), where a set of prompt embeddings is trained in order to improve performance in some specific task. In our case, a shared token $e$ could be viewed as an "instruction" saying what an LLM should do with the upcoming embeddings ($m$-tokens) – decode different pieces of information for different positions.
If the $m$-token is shared, then training and prediction scheme resembles some of the speculative decoding approaches Xia et al. (2023), where a number of special [mask] tokens are appended at the end of the sequence and the prediction for all them is then done in parallel. For all other experiments, unless stated otherwise, we use scheme with sharing $m$ token between texts and random seeds and $e$ token being unique for each text-seed pair.
@@
We already see that similar to autoregressive mode (Kuratov et al., 2025), LLMs can generate fairly long sequences in just one forward pass. To characterize this capability and understand how it scales with model size and changes depending on the nature of the text, we run the optimization process for text prefixes of the predefined lengths [4, 5, 8, 11, 16, 22, 32, 45, 64, 90, 128, 181, 256, 362, 512, 724, 1024, 1448]. We report the maximum values of $C_{tokens}$ and $H_{LM}$ which correspond to the longest prefix for which at least $0.99$ token-level accuracy is achieved – we treat such sequences as successfully reconstructed. In addition to a scheme with a shared $m$ token, we also run a scheme with $m$ not shared, to eliminate the effect of the insufficient number of random initializations. While our results in Section 4, suggest that $m$, can in principle, be shared without any quality drop, we also note that the optimization process is highly sensitive to initialization, especially when the proto-tokens are shared. The results are presented in Table 3 with the best results across datasets presented in Figure 1.
@@
Larger models in Llama the family show greater reconstruction capabilities than the smaller ones, while the situation with Pythia models is less obvious, with all the models showing approximately the same performance. Llama 1B model is also able to reconstruct almost three times larger sequences compared to Pythia model of the same size.
@@
Figure 3: Maximum language information ($H_{LM}$ for a maximum text prefix that is accurately reconstructed) compressed for different models and datasets. In the left plot, a single [mem] token is used in the autoregressive setting, and in the non-autoregressive one, $m$ proto-token is shared between all texts within each model. In the right plot, two [mem] tokens are used and $m$ proto-tokens are not shared. Each small point on the plots represents a single text, larger points indicate the average within each (model, dataset) pair.
@@
The natural text source (unseen, seen or generated) does not seem to have any systematic influence on the quality of reconstruction in terms of the number of tokens, while for unnatural random texts the generation capacity is significantly worse. This suggests that "proto-tokens" do not "store" tokens directly, but encode some more high-level representations, using language modeling capabilities of LLM. However, we also can not say that the compressibility of the text is determined by its likelihood under the sequential language model. In fact, we observe the opposite trend – lower total information content $H_{LM}$ is compressed for less information-dense texts, such as generated by the LLM itself. This difference is highlighted in Figure 3, where the amount of the information contained in trainable tokens is compared to autoregressive setup. The performance for unnatural texts is very similar and sometimes even identical, while for natural texts, the difference in capacity can be up to five times lower. However, more often the difference is just two-fold, suggesting that autoregressive decoding approximately doubles the "effective" information density in natural text.
@@
Figure 4: Reconstruction throughput for autoregressive and non-autoregressive setups. For each model-dataset pair, the throughput equals to a maximum losslessly compressible length divided by the reconstruction time.
@@
Although less information-dense, our one-forward method achieves significantly higher decoding throughput in the context of text reconstruction – outperforming its autoregressive counterpart by a factor of 279 on average (Figure 4). This dramatic difference is due to the number of forward passes. While an obvious downstream task is still to be found, such speed could matter in many settings where fast decoding is particularly important.
While we do not introduce the method as a practical way of compressing or generating texts, but rather as a demonstration of interesting phenomenon, we still measure the training time across models and text lengths to demonstrate the full picture. Training time (Table 4) scales roughly linearly with sequence length, with around $10$ seconds for length $32$ and around $200$ seconds for length $512$.
Model | N $32$ $64$ $128$ $256$ $512$
Pythia-160M 66 $25$ $68$ - -
Pythia-410M $10$ 2121 $106$ - -
Pythia-1.4B $10$ $66$ $129$ - -
Llama-3.2-1B $11$ $15$ $27$ $87$ -
Llama-3.2-3B $14$ $18$ $28$ $78$ $261$
Llama-3.1-8B $16$ 2020 $29$ $60$ $215$
Table 4: Proto-token training time in seconds for different models and sequence lengths averaged over datasets.
@@
We start by measuring the distances between three types of proto-token embedding pairs: 1) corresponding to the same generated sequence, but different random seeds, 2) corresponding to the different texts but generated from the same context, 3) corresponding to the different texts generated from different contexts. As shown in Figure 6, the same-text solutions are almost always located closer to each other than different-texts solutions, which suggests locality in the learned representations. At the same time, same-context solutions are noticeably closer to each other than different-context ones. This may indicate that the encoded information at least partially reflects the potential context of the text. However, we should be careful to account for the texts generated from the same context being more similar in general.
To do that, we measure pairwise distances between generated texts, and examine whether the distance between learned proto-token embeddings differ for a fixed distance between the texts. We use token-level measure of text similarity and semantic-level measure (see Section 3). For both measures, (Figure 5) we observe that, given the same distances between texts, the proto-token embeddings are on average closer when the texts originate from the same context. We conclude that learned proto-tokens contain information beyond the information about the target sequence itself – they somehow partially describe the potential context of the sequence. However, we should note that, the effect of having the same context on the distance between proto-tokens is small, and the distributions for same-context distances and different-context distances heavily overlap. Our results suggest that proto-tokens still mostly contain information about the text itself with only a fraction of the information about the context preserved.
We also conducted a preliminary experiment on accessing the information contained in proro-tokens without first decoding them into text. We took 50 128-token context sequences from PG-19 dataset, generated 256-token continuations with Llama3.2-1B model and trained $(e,m)$ pairs only for the first 128 tokens of the model continuations. Then we started the autoregressive generation of the same model from different combinations of proto-tokens and $\langle\text{BOS}\rangle$-token and visually compared the contents of the resulting token sequences with both contexts and model-continuations. In all cases, the resulting sequences either contain meaningless token-combinations or meaningful texts that are not related to either context or continuation. We conclude that the information from proto-tokens could not be accessed without decoding at least when they are used directly as an autoregressive generation context.
Proto-tokens embedding space structure
Kuratov et al. (2025) raised the following concern about the structure of the solution space in the autoregressive setup. Even though the same-text token embeddings are on average closer to each other than different-text token embeddings, they seem to be disconnected – a linear interpolation between two solutions does not yield a valid reconstruction. This could mean that the potential encoding to this space could be problematic as the same object could be mapped to disconnected regions. We find that in our non-autoregressive case, the linear interpolation between same-text solutions also does not produce a solution (Figure 7).
Figure 7: Pairwise interpolation accuracies between $10$ solutions for 55 texts ($5\times 10\times 9/2$ pairs in total).